OpenStudy (anonymous):

Tim is twice as old as his son. In six years, Tim's age will be three times what his son's age was six years ago. How old is Tim's son now? @mathstudent55

OpenStudy (jdoe0001):

well, let's see tim is twice as old as his son tim = son x 2 let's call t = tim s = son t = s * 2 => t = 2s in 6 years, tim will be 3 times older than his son 6 years ago if we fast-forward 6years, and we ask ourselves, how old was his son 6 years ago? well, we fast-forwarded 6 years, 6 years ago will be TODAY and today his age is "2s" tim's age will be today's age + 6 t + 6 = 3(son's age today) t + 6 = 3(2s) we know t = 2s so we can say (2s) + 6 = 3(2s) solve for "s" to get the Son's age today